9)A New York City daily newspaper called “Manhattan Today” charges an annual subscription fee of $108. Customers prepay their subscriptions and receive 260 issues over the year. To attract more subscribers, the company offered new subscribers the ability to pay $110 for an annual subscription that also would include a coupon to receive a 40% discount on a one-hour ride through Central Park in a horse-drawn carriage. The list price of a carriage ride is $100 per hour. The company estimates that approximately 30% of the coupons will be redeemed.

Required:
1. How much revenue should Manhattan Today recognize upon receipt of the $110 subscription price? (1Revenue_______)
2. How many performance obligations exist in this contract? (Number of performance obligations _________)


3. Prepare the journal entry to recognize sale of 14 new subscriptions, clearly identifying the revenue or deferred revenue associated with each performance obligation.
Answer #1
1
Revenue = 0
2
Number of performance obligations = 2
3
Debit Credit
Cash 1540 =110*14
      Deferred revenue—subscription 1386 =1540*90%
      Deferred revenue—discount coupon 154 =1540*10%
Workings:
Stand­alone selling price of coupon 12 =100*40%*30%
Stand­alone selling price of subscription 108
Total of stand­alone prices 120
Percentage allocated to—subscription 90% =108/120
percentage allocated to—discount coupon 10% =12/120